(1 point) Hannibal opens a savings account on January 1, 1984 with a deposit of 330 dollars, and continues to make deposits of the same amount at the beginning of each month until January 1, 1990, when he makes the final deposit. If the account pays a nominal rate of interest of 5.9 percent convertible monthly, how much is in the account on January 1, 1998?
Answer #1

х New Microsoft Office Excel Worksheet Microsoft Excel Cn Seed Data Home Insert Page Layout Formulas Review View E AutoSum Cuх New Microsoft Office Excel Worksheet - Microsoft Excel Page Layout Data View Home Insert Formulas Review Z AutoSum Cut Wrap

future value =Using future value function in MS excel fv(rate,nper,pmt,pv,type) rate = .4916% nper = 12*6 =72 pmt = -330 pv =0 type =1 FV(0.4916%,72,-330,0,1) $28,565.11
value of investment Jan 1 1998 =Using future value function in MS excel fv(rate,nper,pmt,pv,type) rate = .4916% nper = 12*8 =96 pmt = 0 pv = -28565.11 type =1 FV(0.4916%,96,0,-28565.11,1) $45,739.66